# Efficient Massive Black Hole Binary parameter estimation for LISA using Sequential Neural Likelihood

### v2

@@ 1 Introduction
The first direct detection of gravitational waves (GWs) from a binary black hole merger by the Laser Interferometer Gravitational-Wave Observatory (LIGO) in 2015 [1] has kick-started a new era in astronomy, where GWs constitute a new cosmic messenger that is providing invaluable information for astrophysics, cosmology, and even fundamental physics. GW astronomy is advancing fast, as shown by the number of detections (more than 90 published so far) that have been presented and studied by the LIGO-Virgo collaboration first [1, 2, 3] and, more recently, by the LIGO-Virgo-KAGRA (LVK) collaboration [4], including the first multi-messenger observation involving GWs [5, 6]. All these observations, in the high-frequency GW band, have already provided revolutionary discoveries with impact in astrophysics, cosmology and even fundamental physics. In parallel, in the very-low-frequency GW band (around the nanohertz), pulsar timing array collaborations have recently reported evidence for a GW background [7, 8, 9, 10]. There are also important developments towards the measurement of the cosmic microwave background polarization in order to put constraints on primordial GWs (see, e.g. [11]).
On the other hand, due to the presence of various sources of noise, mainly seismic and (Newtonian) gravity gradient noises, the frequency range to which current ground-based detectors are sensitive is limited to 101010\,10Hz to 101010\,10kHz [12]. Third generation ground-based detectors like Einstein Telescope [13] and Cosmic Explorer [14] can push the low limit further up to 111\,1Hz, while atom-interferometric detectors [15, 16] have been proposed to push it up to the decihertz. Below this, the best option we currently have is to go to space. The Laser Interferometer Space Antenna (LISA) [17] is a space mission led by the European Space Agency (ESA), in collaboration with the US National Aeronautics and Space Administration (NASA), that will survey the observable universe using low-frequency GWs.
@@
In this way, LISA will be sensitive to GW sources in the millihertz band, where we expect to detect GW signals from: (i) The coalescence of massive black hole (MBH) binaries (MBHBs), with total mass in the range $10^{5}-10^{8}\,M_{\odot}$; (ii) the capture and inspiral of a stellar-mass compact object into a MBH, known as an Extreme-Mass-Ratio Inspirals (EMRIs), and similar systems known as Intermediate-Mass-Ratio Inspirals (IMRIs) involving intermediate-mass black holes (IMBHs) [^1] By definition, IMBHs have masses in the range $10^{2}-10^{5}\,M_{\odot}$. Then, we can distinguish between light IMRIs, consisting in the inspiral of an stellar-origin black hole (SOBH) into an IMBH, and heavy IMRIs, consisting of an IMBH inspiralling into a MBH.; (iii) millions of ultra-compact galactic binaries (GBs), whose population is dominated by white dwarf binaries; (iv) Stellar-origin black hole (SOBH) binaries (like GW150914 [1]) in the stage when their periods are in the 111\,1hour range; (v) stochastic GW backgrounds of cosmological origin, associated to high-energy processes at the TeV energy scale, and; (vi) given that LISA is expected to open the low-frequency band, we may find unforeseen GW sources for which we should be prepared.
@@
The current state of the art in LISA data analysis consists of using Bayesian sampling methods to draw samples from the posterior distribution $p(\bm{\theta}\mid{\bf x})$, which is in turn proportional to a likelihood function $p({\bf x}\mid\bm{\theta})$ (here $\bm{\theta}$ denotes the GW model parameters and ${\bf x}$ the detector data). Among the most popular ones we have Markov Chain Monte Carlo (MCMC) sampling methods [39, 40, 41] (see [42, 43, 44, 29, 45] for recent applications to LISA) and nested sampling methods [46, 47] (for recent applications to LISA, see [48, 49, 50]). These methods, while generally reliable, require the repeated evaluation of the likelihood, which in turn demands a large number of waveform simulations during the inference process. Despite the fact that they can be optimized for the problem in question, this ultimately results in high computational costs. Moreover, the likelihood function typically used in GW data analysis comes from the assumption that the noise is stationary and Gaussian, so that it is completely described by the power spectral density (PSD), for which it is also assumed that we have an accurate model. However, this does not reflect reality, indicating that likelihood-based methods will need to either acknowledge the systematic errors introduced, or to try to take them into account by adding corrections to the likelihood – further driving up the computational cost.
The main goal of this work is to propose, as an alternative approach to likelihood-based methods, the use of simulation-based inference (SBI) techniques [51]. SBI leverages the recent developments in Machine Learning to obtain the posterior distribution using only the information contained within simulated data. All assumptions about the data – aside from the prior – are thus built into the simulator itself. As such, there is no need to define a likelihood function, making SBI applicable to problems where it is not possible to define a likelihood function or where it is numerically expensive to compute it. Moreover, SBI techniques have already seen remarkable success when applied to GW source parameter estimation in the case of ground-based detectors. Most notably, DINGO [52, 53, 54] is able to do amortized inference: after an initial training, a full inference run on previously unseen data with state-of-the-art accuracy can be performed in seconds to minutes, instead of hours to months.
@@ 2 Likelihood-based inference basics
The sum in this inner product is over all the detector channels, and $[S_{n}(f)]^{ij}$ denotes the noise power spectral density matrix correlated between detector channels $i$ and $j$. In the case of ground-based detectors, it is common to assume that the noises at each detector are uncorrelated, and therefore $[S_{n}(f)]^{ij}=0$ for $i\neq j$. In LISA, however, the noises at each spacecraft are correlated, which translates into correlations between the TDI observables $X$, $Y$, and $Z$. Instead, we can construct TDI combinations $(A,E,T)$ that are uncorrelated, which simplifies the likelihood computation. Furthermore, the sensitivity to GWs of the TDI channel $T$ is heavily suppressed at low frequencies [60, 23].
@@
Given the likelihood function defined in Eq. (2.3), one can sample from the posterior distribution using Eq. (2.1). The evidence, $p({\bf x})$, is usually numerically intractable and hence it is usually ignored and normalization of the posterior probability distribution is performed after sampling. The evidence becomes relevant when the number of sources present in the data is unknown, or when comparing between different waveform models, but those scenarios lie outside of the scope of this work.
@@ 4 Simulation pipeline
In order to train our model, large amounts of simulated MBHB coalescence data are needed. We have created a data generation pipeline that can rapidly ‘mass-produce’ LDC-like datasets with arbitrary parameters. Speed of simulation was prioritized over data fidelity and realism, in order to be able to rapidly develop, prototype and test our algorithms. In this sense, our choice of waveform model is the frequency-domain IMRPhenomD [56, 57] model, which describe the inspiral, merger and ringdown phases of non-precessing (aligned-spin) binary BH coalescences. Similarly, the preprocessing of the data makes several assumptions — detailed in what follows — that, for the sake of speed, may not be completely realistic. Because of this, it is expected that our inference will not give satisfactory results when shown real data, but in the same way as other methods that use the same type of waveform model. Nevertheless, it is always possible to replace our simulation pipeline with a more realistic one, without changing the overall scheme of the method, and train using data generated from it to obtain more robust results. This is going to be part of future improvements of the method. A summary of the procedure is given in Algorithm 1, with more details given in the following subsections.
@@ 4.1 Choice of initial priors
$\chi_{1}$ Projected spin [^6] IMRPhenomD only supports BH spins perpendicular to the orbital plane. The projected spin $\chi_{i}$ is the component of the spin projected on that axis. If the dimensionless spin ranges from 00 to $1$, $\chi_{i}$ can range from $-1$ to $1$. of BH 1 0.26863190922667673 $\mathcal{U}(\chi_{1};-1,1)$
@@ 4.2 Waveform generation and noise
For every set of MBHB parameters $\bm{\theta}$, a MBHB simulation is then run using the lisabeta code (see [79, 44]). The waveform model supplied by this code is an accelerated version of the IMRPhenomD model [56, 57], using the formalism described in Ref. [79] to apply the LISA instrumental response in the frequency domain and directly output TDI observables. The comparatively slower time-domain calculation is thus avoided. The outputs of the code are the second-generation TDI $(A,E,T)$ channels [24, 23] in frequency domain – though the $T$ channel was promptly discarded, as the amount of GW information it carries is suppressed at low frequencies [60, 23]. Finally, Keplerian LISA orbits are used in the computation of the waveform response.
@@ 6.1 Noise-free case
We begin by showing the results in the absence of noise. Fig. 4 shows the parameters recovered by Eryn and Fig. 5 shows the results obtained with SNL-PCA. In both cases, for the intrinsic parameters — as shown in the top right sections of each — the recovered posteriors do have a maximum around the true parameter values, so they are correctly recovered by both algorithms. The most notable difference between both posteriors is in the $m_{1}$$m_{2}$ 2D posteriors, which have different shapes. This is because SNL-PCA has not been able to recover the chirp mass $\mathcal{M}_{c}$ with the same accuracy as Eryn, but in turn, it constrained the mass ratio $q$ remarkably better. Thus, the elongated shape of the $m_{1}$-$m_{2}$ posterior in Fig. 4 follows a line of constant $\mathcal{M}_{c}$, while the corresponding panel in Fig. 5 is more distributed along the direction of constant $q$. The 1D projections for each intrinsic parameter are ultimately of similar width, with SNL-PCA being only slightly narrower than Eryn.
@@
As for SNL-AE, Fig. 6 shows the resulting posterior. We can see that our algorithm is able to successfully extract the intrinsic parameters of the injected MBHB source. The results in these parameters show posteriors noticeably broader than Eryn (Fig. 4), but otherwise they are qualitatively identical.
@@ 6.2 Noisy case
The presentation of the results with noise follows the same line as in the noise-free case. Fig. 7 shows the parameters recovered by Eryn, while Fig. 8 shows the SNL-PCA equivalent. Comparing them, we can see qualitatively similar results for the intrinsic parameters, where the posteriors have similar shapes and are centered around the true values, but the SNL-PCA posteriors are slightly broader due to lower fidelity in recovering the chirp mass.
@@
Finally, Fig. 9 shows the posterior obtained for SNL-AE. Extraction of the intrinsic parameters from the posteriors becomes a lot more difficult for SNL-AE. For the masses, while the 1D marginalized posteriors may be off from their true value, in the $m_{1}$-$m_{2}$ projection we can see that the estimated masses are very well localized on a curve of constant chirp mass: the chirp mass is well recovered, in contrast with the mass ratio. A similar behaviour is present in the spins, where the mode of $\chi_{1}$ is only somewhat near the true value, but the projection in the $\chi_{1}$-$\chi_{2}$ plane shows that the spins are rather well-determined in some combination of $\chi_{1}$ and $\chi_{2}$ – likely related to the leading-order Post-Newtonian spin parameterization $\chi_{PN}$ that plays a role in the IMRPhenomD waveform model [56, 57]. It should be noted that the upper right corner of Fig. 9 looks qualitatively similar to Fig. 8 of Ref. [42]. This figure shows the posterior for the intrinsic parameters for a pre-merger MBHB signal – that is, including only the inspiral part of the waveform. The fact that our results look similar to that scenario reinforces the idea that the degradation in our SNL-AE posteriors can be explained by the fact that our autoencoder is failing to reconstruct the merger, as explained in Sec. 5.2 (and shown explicitly in Figs. 3(c) and 3(d)).

### v3

@@ 1 Introduction
The first direct detection of gravitational waves (GWs) from a binary black hole merger by the Laser Interferometer Gravitational-Wave Observatory (LIGO) in 2015 [1] has kick-started a new era in astronomy, where GWs constitute a new cosmic messenger that is providing invaluable information for astrophysics, cosmology, and even fundamental physics. GW astronomy is advancing fast, as shown by the number of detections (more than 90 published so far) that have been achieved by the second generation of ground-based interferometric detectors [2, 3, 4] and presented by the LIGO-Virgo collaboration first [1, 5, 6] and, more recently, by the LIGO-Virgo-KAGRA (LVK) collaboration [7], including the first multi-messenger observation involving GWs [8, 9]. Additional events outside the LVK collaboration have been reported in [10, 11, 12, 13, 14, 15]. All these observations, in the high-frequency GW band, have already provided revolutionary discoveries with impact in astrophysics, cosmology and even fundamental physics. In parallel, in the very-low-frequency GW band (around the nanohertz), pulsar timing array collaborations have recently reported evidence for a GW background [16, 17, 18, 19]. There are also important developments towards the measurement of the cosmic microwave background polarization in order to put constraints on primordial GWs (see, e.g. [20]).
On the other hand, due to the presence of various sources of noise, the sensitivity of ground-based detectors is limited to the frequency range between 101010\,10Hz and 101010\,10kHz [21], mainly due to seismic and (Newtonian) gravity gradient noises in the low end, and to quantum noise in the high end. Third generation ground-based detectors like Einstein Telescope [22] and Cosmic Explorer [23] can push the low limit further up to 111\,1Hz, while atom-interferometric detectors [24, 25] have been proposed to push it up to the decihertz. Below this, the best option we currently have is to go to space. The Laser Interferometer Space Antenna (LISA) [26] is a space mission led by the European Space Agency (ESA), in collaboration with the US National Aeronautics and Space Administration (NASA), that will survey the observable universe using low-frequency GWs.
@@
In this way, LISA will be sensitive to GW sources in the millihertz band, where we expect to detect GW signals from (see, e.g. [26, 27]): (i) The coalescence of massive black hole (MBH) binaries (MBHBs), with total mass in the range $10^{5}-10^{8}\,M_{\odot}$; (ii) the capture and inspiral of a stellar-mass compact object into a MBH, known as an Extreme-Mass-Ratio Inspirals (EMRIs), and similar systems known as Intermediate-Mass-Ratio Inspirals (IMRIs) involving intermediate-mass black holes (IMBHs) [^1] By definition, IMBHs have masses in the range $10^{2}-10^{5}\,M_{\odot}$. Then, we can distinguish between light IMRIs, consisting in the inspiral of an stellar-origin black hole (SOBH) into an IMBH, and heavy IMRIs, consisting of an IMBH inspiralling into a MBH.; (iii) millions of ultra-compact galactic binaries (GBs), whose population is dominated by white dwarf binaries; (iv) Stellar-origin black hole (SOBH) binaries (like GW150914 [1]) in the stage when their periods are in the 111\,1hour range; (v) stochastic GW backgrounds of cosmological origin, associated to high-energy processes at the TeV energy scale, and; (vi) given that LISA is expected to open the low-frequency band, we may find unforeseen GW sources for which we should be prepared.
@@
The current state of the art in LISA data analysis consists of using Bayesian sampling methods to draw samples from the posterior distribution $p(\bm{\theta}\mid{\bf x})$, which is in turn proportional to a likelihood function $p({\bf x}\mid\bm{\theta})$ (here $\bm{\theta}$ denotes the GW model parameters and ${\bf x}$ the detector data). Among the most popular ones we have Markov Chain Monte Carlo (MCMC) sampling methods [48, 49, 50] (see [51, 52, 53, 38, 54] for recent applications to LISA) and nested sampling methods [55, 56] (for recent applications to LISA, see [57, 58, 59] [^2] Refs. [58, 59] adapt PyCBC [60] and bilby [61] respectively for their use in LISA data analysis. They perform nested sampling by using their interfaces to the dynesty [62, 63] sampler.), but there are other options (see, e.g. [64]). These methods, while generally reliable, require the repeated evaluation of the likelihood, which in turn demands a large number of waveform simulations during the inference process. Despite the fact that they can be optimized for the problem in question, this ultimately results in high computational costs. Moreover, the likelihood function typically used in GW data analysis comes from the assumption that the noise is (wide-sense) stationary and (multi-variate) Gaussian, so that it is completely described by the power spectral density (PSD), for which it is also assumed that we have an accurate model. However, this does not reflect reality, indicating that likelihood-based methods will need to either acknowledge the systematic errors introduced, or to try to take them into account by adding corrections to the likelihood – further driving up the computational cost.
The main goal of this work is to propose, as an alternative approach to likelihood-based methods, the use of simulation-based inference (SBI) techniques [65]. SBI leverages the recent developments in Machine Learning to obtain the posterior distribution using only the information contained within simulated data. All assumptions about the data – aside from the prior – are thus built into the simulator itself. As such, there is no need to define a likelihood function, making SBI applicable to problems where it is not possible to define a likelihood function or where it is numerically expensive to compute it. Moreover, SBI techniques have already seen remarkable success when applied to GW source parameter estimation in the case of ground-based detectors (see, e.g. [66, 67]). Most notably, DINGO [68, 69, 70] is able to do amortized inference: after an initial training, a full inference run on previously unseen data with state-of-the-art accuracy can be performed in seconds to minutes, instead of hours to months.
@@ 2 Likelihood-based inference basics
The sum in this inner product is over all the detector channels, and $[S_{n}(f)]^{ij}$ denotes the noise power spectral density matrix correlated between detector channels $i$ and $j$. In the case of a network of ground-based detectors situated at different locations, it is common to assume that the noises at each detector are uncorrelated [^4] For the Einstein Telescope, where multiple detectors are planned to be in the same location, this is no longer the case., and therefore $[S_{n}(f)]^{ij}=0$ for $i\neq j$. In LISA, however, the noises at each spacecraft are correlated, which translates into correlations between the TDI observables $X$, $Y$, and $Z$. Instead, we can construct TDI combinations $(A,E,T)$ that are uncorrelated, which simplifies the likelihood computation. Furthermore, the sensitivity to GWs of the TDI channel $T$ is heavily suppressed at low frequencies [76, 32]. In the case of MBHB signals, it is safe to ignore this channel, as we can see from the frequency range covered by MBHBs that is discussed inside Appendix A, in particular in Figure 12.
@@
Given the likelihood function defined in Eq. (2.3), one can sample from the posterior distribution using Eq. (2.1). The evidence, $p({\bf x})$, is usually numerically intractable and hence it is usually ignored and normalization of the posterior probability distribution is performed after sampling [^5] To be more specific, MCMC sampling obtains samples of the (unnormalized) posterior, and by normalizing them one can also obtain an estimate of the evidence. The main goal of nested sampling is to compute the evidence, but one also ends with a set of posterior samples as a byproduct.. The evidence becomes relevant when the number of sources present in the data is unknown, or when comparing between different waveform models, but those scenarios lie outside of the scope of this work.
@@ 4 Simulation pipeline
In order to train our model, large amounts of simulated MBHB coalescence data are needed. We have created a data generation pipeline that can rapidly ‘mass-produce’ LDC-like datasets with arbitrary parameters. Speed of simulation was prioritized over data fidelity and realism, in order to be able to rapidly develop, prototype and test our algorithms. Our choice of waveform model is the frequency-domain IMRPhenomD [72, 73] model, which describes the $\ell=|m|=2$ spin-weighted spherical-harmonic modes of the GW signal emitted during the inspiral, merger and ringdown phases of non-precessing (aligned-spin) binary BH coalescences. The intrinsic parameters of the IMRPhenomD waveform model are: (i) The masses of the BHs, $m_{1}$ and $m_{2}$; and (ii) the projected spins [^9] IMRPhenomD only supports BH spins perpendicular to the orbital plane. The projected spin $\chi_{i}$ is the component of the spin projected on that axis. If the dimensionless spin ranges from 00 to $1$, $\chi_{i}$ can range from $-1$ to $1$. in the direction of the orbital angular momentum, $\chi_{1}$ and $\chi_{2}$. There are some combinations of these parameters that have interest because the GW signal is specially sensitive to them. In the case of masses we have the chirp mass
$\mathcal{M}_{c}=\frac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}=\frac{q^{3/5}}{(1+q)^{1/5}}\,m_{1}\,,$ (4.1)
where $q=m_{2}/m_{1}$ is the mass ratio. In the case of spins we have the effective spin parameter
$\chi_{\mathrm{eff}}=\frac{m_{1}\chi_{1}+m_{2}\chi_{2}}{M}=\frac{\chi_{1}+q\chi_{2}}{1+q}\,,$ (4.2)
where $M=m_{1}+m_{2}$ is the total mass of the binary BH. On the other hand, the extrinsic parameters of the IMRPhenomD model are: (i) the luminosity distance, $D_{L}$; (ii) the longitude and latitude of the binary BH sky location in the LISA frame [53], $\lambda_{L}$ and $\beta_{L}$ respectively; (iii) the coalescence time, $t_{c}$; (iv) the GW phase at coalescence time, $\phi_{c}$; (v) the inclination angle between the orbital angular momentum and the line of sight, $\iota$; and (vi) the GW polarization angle, $\psi_{L}$. In total, the waveform model has 11 parameters (see Table 1 for the values of these parameters assumed in this work and the associated prior distributions).
The preprocessing of the data makes several assumptions — detailed in what follows — that, for the sake of speed, may not be completely realistic. Because of this, it is expected that our inference will not give satisfactory results when shown real data, but in the same way as other methods that use the same type of waveform model. Nevertheless, it is always possible to replace our simulation pipeline with a more realistic one, without changing the overall scheme of the method, and train using data generated from it to obtain more robust results. This is going to be part of future improvements of the method. A summary of the procedure is given in Algorithm 1, with more details given in the following subsections.
@@ 4.1 Choice of initial priors
$\chi_{1}$ Projected spin of BH 1 0.26863190922667673 $\mathcal{U}(\chi_{1};-1,1)$
@@ 4.2 Waveform generation and noise
For every set of MBHB parameters $\bm{\theta}$, a MBHB simulation is then run using the lisabeta code (see [94, 53]). The waveform model supplied by this code is an accelerated version of the IMRPhenomD model [72, 73], using the formalism described in Ref. [94] to apply the LISA instrumental response in the frequency domain and directly output TDI observables. The comparatively slower time-domain calculation is thus avoided. The outputs of the code are the second-generation TDI $(A,E,T)$ channels [33, 32] in frequency domain – though the $T$ channel was promptly discarded, as the amount of GW information it carries is suppressed at low frequencies [76, 32]. In connection to this, in Appendix A we provide information about the maximum frequency of MBHB events (see also Fig. 12), which justifies this approximation. Finally, Keplerian LISA orbits are used in the computation of the waveform response.
@@ 6.1 Noise-free case
(a) Eryn
(b) SNL-PCA
(c) SNL-AE
Figure 6: Noise-free posteriors for the chirp mass $\mathcal{M}_{c}$, the mass ratio $q$ and the spin combination $\chi_{\mathrm{eff}}$.
We begin by showing the results in the absence of noise. Fig. 4 shows the parameters recovered by Eryn and Fig. 5 shows the results obtained with SNL-PCA. In both cases, for the intrinsic parameters — as shown in the top right sections of each — the recovered posteriors do have a maximum around the true parameter values, so they are correctly recovered by both algorithms. The 1D projections for each intrinsic parameter are ultimately of similar width, with SNL-PCA being only slightly narrower than Eryn. The most notable difference between both posteriors is in the $m_{1}$$m_{2}$ 2D posteriors, which have different shapes. In order to diagnose this, we show in Fig. 6 the posteriors after reparameterizing to the chirp mass $\mathcal{M}_{c}$, the mass ratio $q=m_{2}/m_{1}$, and the effective spin $\chi_{\mathrm{eff}}$ (see Eqs. (4.1) and (4.2)). Comparing Figs. 6(a) and 6(b), we can see that SNL-PCA has not been able to recover the chirp mass $\mathcal{M}_{c}$ with the same accuracy as Eryn, but in turn, it constrained the mass ratio $q$ remarkably better. Thus, the elongated shape of the $m_{1}$-$m_{2}$ posterior in Fig. 4 approximately follows a line of constant $\mathcal{M}_{c}$, while the corresponding panel in Fig. 5 is more distributed along the direction of constant $q$. The effective spin parameter $\chi_{\mathrm{eff}}$, as defined in Eq. (4.2), depends only on $q$ and the individual spins $\chi_{1}$ and $\chi_{2}$ — the fact that SNL-PCA is able to constrain these three parameters better than Eryn explains the narrower effective spin posteriors.
@@
As for SNL-AE, Fig. 7 shows the resulting posterior. We can see that our algorithm is able to successfully extract the intrinsic parameters of the injected MBHB source. The results in these parameters show posteriors noticeably broader than Eryn (Fig. 4), but otherwise they are qualitatively identical. The broadening of the posterior for the intrinsic parameters translates into much wider posteriors for the intrinsic parameter combinations in Fig. 6(c).
@@ 6.2 Noisy case
(a) Eryn
(b) SNL-PCA
(c) SNL-AE
Figure 10: Noisy posteriors for the chirp mass $\mathcal{M}_{c}$, the mass ratio $q$ and the spin combination $\chi_{\mathrm{eff}}$.
The presentation of the results with noise follows the same line as in the noise-free case. Fig. 8 shows the parameters recovered by Eryn, while Fig. 9 shows the SNL-PCA equivalent. Comparing them, we can see qualitatively similar results for the intrinsic parameters, where the posteriors have similar shapes and are centered around the true values, but the SNL-PCA posteriors are slightly broader due to lower fidelity in recovering the chirp mass. Fig. 10 shows the intrinsic posteriors of the chirp mass $\mathcal{M}_{c}$, the mass ratio $q$ and the effective spin $\chi_{\mathrm{eff}}$. Here, we see that the SNL-PCA posteriors of Fig. 10(b) remain compatible with the truth — although they are slightly broader than Eryn’s (Fig. 10(a)), and the correlation between parameter pairs is lost.
@@
Finally, Fig. 11 shows the posterior obtained for SNL-AE. Extraction of the intrinsic parameters from the posteriors becomes a lot more difficult for SNL-AE. For the masses, while the 1D marginalized posteriors may be off from their true value, in the $m_{1}$-$m_{2}$ projection we can see that the estimated masses are very well localized on a curve of constant chirp mass: the chirp mass is well recovered, in contrast with the mass ratio. Indeed, Fig. 10(c) shows that the one-dimensional posterior for the chirp mass $\mathcal{M}_{c}$ is of a similar width and shape to the one recovered by SNL-PCA (Fig. 10(b)), while the mass ratio $q$ remains badly constrained. The individual spins, $\chi_{1}$ and $\chi_{2}$, are also poorly recovered — as can be seen in Fig. 11, while the mode of $\chi_{1}$ is somewhat near the true value, the marginal posteriors for both parameters occupy nearly the entire range of possible values. However, the projection in the $\chi_{1}$-$\chi_{2}$ plane shows that the spins are rather well-determined in a combination of $\chi_{1}$ and $\chi_{2}$ — just like the chirp mass was compared to the individual mass posteriors. The $\chi_{\mathrm{eff}}$ posterior, shown in Fig. 10(c), is significantly narrower than either of the individual spins, but the shape remains irregularly-shaped, biased from the true value, and much broader than both methods. It should be noted that the upper right corner of Fig. 11 looks qualitatively similar to Fig. 8 of Ref. [51]. This figure shows the posterior for the intrinsic parameters for a pre-merger MBHB signal – that is, including only the inspiral part of the waveform. The fact that our results look similar to that scenario reinforces the idea that the degradation in our SNL-AE posteriors can be explained by the fact that our autoencoder is failing to reconstruct the merger, as explained in Sec. 5.2 (and shown explicitly in Figs. 3(c) and 3(d)).